# Recursive Quantum Relaxation for Combinatorial Optimization Problems

### v2

@@ 3.2 Theoretical Analysis
The proposed method, RQRAO, combines the RQAOA [3] and QRAO [11]. These existing methods work well experimentally, but there are some theoretical ambiguities: (i) Why does the recursive approach improve the cut weight performance in the RQAOA? (ii) What is the meaning of maximizing the expectation value of the QRAC Hamiltonian, and what is the theoretical background of Pauli rounding in QRAO? To answer these questions, we prove the following theorem:
@@
${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}\in\{0,1\}^{|V|}}\mathcal{P}_{m}({\bf b};\rho^{*}),$ (19)
where $\mathcal{P}_{m}({\bf b};\rho)$ is the probability of obtaining ${\bf b}$ by $(m,1)$-random magic measurements and is proportional to the fidelity between $\rho$ and the $(m,1)$-magic state embedding $\mu({\bf b})$.
From Theorem 1, it is clarified that both the RQAOA and QRAO, let alone RQRAO, are models searching for the highest probability sample of the probability distribution that maximizes the expectation value of the cut weight through the quantum state as the hidden variable. It will be proved later that maximizing the expectation value of the QRAC Hamiltonian, let alone that of the Ising-type Hamiltonian, can be seen as carrying out Eq. (18). In addition, it will also be proved later that both the recursive approach in the RQAOA and Pauli rounding in the QRAO are just methods that approximate methods of Eq. (19).
Of course, Theorem 1 does not say the MAX-CUT problem is efficiently solvable because both maximizing the expected cut weight and obtaining the highest probability sample is, in general, intractable when ${\bf b}$ lives in an extremely high dimensional space, $\{0,1\}^{|V|}$ for $|V|\gg 1$. However, we can say that it is natural to take the ${\bf b}$ with the highest probability because, after solving Eq. (18), which maximizes the expected cut weight, the probability for larger $\texttt{CW}({\bf b})$ is expected to be large. It is therefore reasonable calculate Eqs. (18) and (19) to obtain a better cut weight.
In Sec. 3.2.1, Eq. (18) is proved. Then, Eq. (19) is proved in Sec. 3.2.2. All the proofs of the following theorems and corollaries are deferred to Supplementary Information C.
@@
The objectives of the RQAOA, QRAO, and RQRAO are to maximize ${\rm tr}(H_{m}\rho)$, where $m=1$ for the RQAOA whereas $m\in\{1,2,3\}$ for QRAO and RQRAO. Hence, to prove Eq. (18), we need to show that maximizing ${\rm tr}(H_{m}\rho)$ is equivalent to maximizing $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}$. To show this, let us start with the following theorem and corollary, which are proved for the first time in this paper.
@@
According to Eq. (22), the expectation value of the $(m,1)$-QRAC Hamiltonian is proportional to the expectation value of the cut weight, where the expectation is taken by the bit strings obtained by $(m,1)$-random magic measurements. Hence, maximizing the expectation value of the $(m,1)$-QRAC Hamiltonian can be regarded as maximizing the expected cut weight. Note that the lower bounds of $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}$ for $w_{jk}=1$ when ${\rm tr}(H_{m}\rho)\geq\texttt{CW}({\bf b}^{*})$ have been given in [11, Theorem 2 and Supplementary Information VII]. They are $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}\geq\frac{5}{9}\texttt{CW}({\bf b}^{*})$ for $m=3$ and $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}\geq\frac{5}{8}\texttt{CW}({\bf b}^{*})$ for $m=2$. These bounds are reproduced by substituting ${\rm tr}(H_{m}\rho)\geq\texttt{CW}({\bf b}^{*})$ and $\sum_{(j,k)\in E}w_{jk}\geq\texttt{CW}({\bf b}^{*})$ into Eq. (22).
@@
First, Eq. (19) is rewritten in the form of the conditional distribution to make it easier to understand. A joint distribution $\mathcal{P}_{m}({\bf b};\rho^{*})$ can be written as the product of the conditional distribution as
$\mathcal{P}_{m}({\bf b};\rho^{*})=\prod^{|V|}_{a=1}\mathcal{P}_{m}(b_{j_{a}}|{\bf b}_{<j_{a-1}};\rho^{*}),$ (24)
where ${\bf b}_{<j_{a-1}}:=\{b_{j_{1}},\ldots,b_{j_{a-1}}\}$. Using this relationship, Eq. (19) can be written as
${b^{\rm R}_{j_{a}}=\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\mathcal{P}_{m}(b_{j_{a}}|{\bf b}^{\rm R}_{<j_{a}};\rho^{*})}$
${\rm for}\ \ \ a=1,2,\ldots,|V|.$ (25)
Because the cut weight is invariant to the bit flip, we can fix one of the $b^{\rm R}_{j}$ to 00 or $1$. In this case, there exists a one-to-one relationship between $\mathcal{P}_{m}({\bf b};\rho^{*})$ and
${\mathcal{P}}_{m}(\mathfrak{p};\rho^{*},T):=\prod^{|V|-1}_{b=1}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{<e_{b}};\rho^{*},T),$ (26)
@@
is a random variable of an edge parity, where “$+$” and “--” indicate positive and negative parities, respectively, $T=\{e_{j_{b}}=(c_{j_{b}},p_{j_{b}})\}^{|V|-1}_{b=1}$ is an arbitrary spanning tree of given graph $G$, and $\mathfrak{p}_{<e_{b}}:=\{\mathfrak{p}_{e_{1}},\ldots,\mathfrak{p}_{e_{b-1}}\}$. Here, the reason for considering the probability of the edge parity on the tree is to exclude inconsistent parities such as a triangle with negative parities. Excluding such an inconsistency does not lose the generality because all possible configurations of solutions can be treated by considering only the probability on the tree. From the view of edge parity, Eq. (19) can be rewritten as
${\mathfrak{p}^{\rm R}_{e_{b}}=\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}^{\rm R}_{<e_{b}};\rho^{*},T^{*})}$
${\rm for}\ \ \ b=1,2,\ldots,|V|-1,$ (28)
where
$T^{*}:=\mathop{\rm argmax}_{T}\max_{\mathfrak{p}\in\{+,-\}^{|V|-1}}{\mathcal{P}}_{m}(\mathfrak{p};\rho^{*},T).$ (29)
In the following, we show that QRAO uses Eq. (25), whereas both the RQAOA and RQRAO use Eq. (28) to obtain the solution. All of them assume a moderate approximation on a probability distribution. The following corollary is frequently used for the proof.
@@
Assume all $b_{j_{a}}$ for $a=1,2,\ldots,|V|$ are independent. Under this assumption, Eq. (24) is approximated as
${\mathcal{P}}^{\rm Pauli}_{m}({\bf b};\rho^{*}):=\prod^{|V|}_{a=1}\mathcal{P}_{m}(b_{j_{a}};\rho^{*}).$ (32)
Substituting Eq. (32) into Eq. (25) gives
$\begin{array}[]{lll}b^{\rm Pauli}_{j_{a}}&:=&\displaystyle{\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\mathcal{P}_{m}(b_{j_{a}};\rho^{*})}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{j_{a}}}}{2\sqrt{m}}\mathcal{E}_{j_{a}}\Bigg{)}}\\[14.22636pt] &=&\left\{\begin{array}[]{ll}0&(\mathcal{E}_{j_{a}}>0)\\ 1&(\mathcal{E}_{j_{a}}<0)\end{array}\right.,\end{array}$ (33)
where the second equality uses Corollary 3. Interestingly, ${\bf b}^{\rm Pauli}$ is equivalent to the result of the Pauli rounding of QRAO [11]. This comes out that the Pauli rounding seeks the variable with the highest probability of ${\mathcal{P}}_{m}({\bf b};\rho)$ under the assumption of Eq. (32).
@@
Assume all $\mathfrak{p}_{e_{b}}$ for $b=1,2,\ldots,|V|-1$ are independent. Under this assumption, Eq. (26) is approximated as
${\mathcal{P}}^{\rm tree}_{m}(\mathfrak{p};\rho^{*},T):=\displaystyle{\prod_{b=1}^{|V|-1}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}};\rho^{*},T)}.\$ (34)
Substituting Eq. (34) into Eq. (28) gives
$\begin{array}[]{lll}\mathfrak{p}^{\rm tree}_{e_{b}=(c,p)}&:=&\displaystyle{\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}};\rho^{*},T^{*})}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{(b_{c},b_{p})\in\{0,1\}^{2}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{c}+b_{p}}}{2m}\mathcal{E}_{cp}\Bigg{)}}\\[14.22636pt] &=&\left\{\begin{array}[]{ll}+&(\mathcal{E}_{cp}>0)\\ -&(\mathcal{E}_{cp}<0)\end{array}\right.,\end{array}$ (35)
where the second equality uses Corollary 3, and
$\begin{array}[]{lll}T^{*}&=&\displaystyle{\mathop{\rm argmax}_{T}\max_{\mathfrak{p}}{\mathcal{P}}^{\rm tree}_{m}(\mathfrak{p};\rho^{*},T)}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{T\ni\{(c,p)\}}\max_{(b_{c},b_{p})\in\{0,1\}^{2}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{c}+b_{p}}}{2m}\mathcal{E}_{cp}\Bigg{)}}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{T\ni\{(c,p)\}}\big{|}\mathcal{E}_{cp}\big{|}}\end{array}$ (36)
is a maximum spanning tree. Computing Eq. (35) can be carried out by a two-step calculation. In the first step, the maximum spanning tree $T^{*}$ is computed with the edge weights $|\mathcal{E}_{cp}|$ for edge $(c,p)$. In the second step, the edge parity is determined by the sign of corresponding $\mathcal{E}_{cp}$. We refer to this rounding method as tree rounding.
RQRAO with $N=1$, which has no recursive steps, is equivalent to using tree rounding. Because both Pauli and tree rounding assume that all random variables are independent, it is expected that QRAO and RQRAO ($N=1$) will yield a similar cut weight.
@@
Both Pauli and tree rounding impose the strong approximation on $\mathcal{P}_{m}$ that all random variables are independent. This is because conditional sampling is in general intractable to carry out. For this reason, let us impose the condition $\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}$ on the optimization problem as the constraint instead of on $\mathcal{P}_{m}$. Of course, these two approaches are not exactly equivalent, but imposing the condition as the constraint of the optimization should be a good approximation to the conditional sampling. The specific steps are given below.
@@
Finally, $\mathfrak{p}^{*}=\{\mathfrak{p}^{*}_{1},\ldots,\mathfrak{p}^{*}_{|V|-1}\}$ is obtained. The corresponding bit string ${\bf b}^{\rm rec}$ can be easily recovered from $\mathfrak{p}^{*}$ by fixing one of the $b^{\rm rec}_{j}$ to 00 or $1$. The condition $\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}$ is indirectly imposed through $\rho^{*}_{b}$ to $\mathcal{P}_{m}$, and hence this rounding method imposes an approximation that is weaker than that of Pauli and tree rounding. We refer to this rounding method as recursive rounding.

### v3

@@ 3.2 Theoretical Analysis
The proposed method, RQRAO, combines the existing RQAOA [3] and QRAO [11]. Although the existing methods work well experimentally, there are interesting theoretical questions: (i) Why does the recursive approach improve the cut weight performance in the RQAOA? (ii) What is the meaning of maximizing the expectation value of the QRAC Hamiltonian, and what is the theoretical background of Pauli rounding in QRAO? To answer these questions, we prove the following theorem:
@@
${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}\in\{0,1\}^{|V|}}\mathcal{P}_{m}({\bf b};\rho^{*})+\mathcal{P}_{m}(\overline{{\bf b}};\rho^{*}),$ (20)
where $\overline{\bf b}$ is the bit-flip of ${\bf b}$, and $\mathcal{P}_{m}({\bf b};\rho)$ is the probability of obtaining ${\bf b}$ by $(m,1)$-random magic measurements whose value is proportional to the fidelity between $\rho$ and the $(m,1)$-magic state embedding $\mu({\bf b})$.
Here, the superscript “R” stands for “rounding” and emphasizes that it represents the classical bits obtained by decoding the information embedded in the qubits. From Theorem 1, it is clarified that both the RQAOA and QRAO, let alone RQRAO, are models searching for the highest probability sample of the probability distribution that maximizes the expectation value of the cut weight through the quantum state as the hidden variable. It will be proved later that maximizing the expectation value of the QRAC Hamiltonian, let alone that of the Ising-type Hamiltonian, can be seen as carrying out Eq. (19). In addition, it will also be proved later that both the recursive approach in the RQAOA and Pauli rounding in the QRAO are just methods that approximate Eq. (20). One of the contributions of this paper is a discovery that QRAO, RQAOA, and RQRAO, which at first glance seem unrelated in terms of their decoding processes, all decode bit strings according to Eq. (20). This allows us to consider a unified framework that encompasses all methods.
Surely, Theorem 1 does not say the MAX-CUT problem is efficiently solvable because both maximizing the expected cut weight and obtaining the highest probability sample is, in general, intractable when ${\bf b}$ lives in an extremely high dimensional space. However, we can say that it is natural to take the ${\bf b}$ with the highest probability because, after solving Eq. (19), which maximizes the expected cut weight, the probability for larger $\texttt{CW}({\bf b})$ is expected to be large. It is therefore reasonable to solve Eqs. (19) and (20) to obtain a better cut weight.
In Sec. 3.2.1, Eq. (19) is proved, while Eq. (20) is proved in Sec. 3.2.2. All the proofs of the theorems and corollaries in the following section are deferred to Supplementary Information C.
@@
The objectives of the RQAOA, QRAO, and RQRAO are to maximize ${\rm tr}(H_{m}\rho)$, where $m=1$ for the RQAOA whereas $m\in\{2,3\}$ for QRAO and RQRAO. Hence, to prove Eq. (19), we need to show that maximizing ${\rm tr}(H_{m}\rho)$ is equivalent to maximizing $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}$. For this, let us start with the following theorems and corollaries, which are given for the first time in this paper.
@@
According to Eq. (23), the expectation value of the $(m,1)$-QRAC Hamiltonian is proportional to the expectation value of the cut weight, where the expectation is over the bit strings obtained from $(m,1)$-random magic measurements. Hence, maximizing the expectation value of the $(m,1)$-QRAC Hamiltonian can be regarded as maximizing the expected cut weight. Note that the lower bounds of $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}$ for $w_{jk}=1$ when ${\rm tr}(H_{m}\rho)\geq\texttt{CW}({\bf b}^{*})$ have been given in [11, Theorem 2 and Supplementary Information VII]. They are $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}\geq\frac{5}{9}\texttt{CW}({\bf b}^{*})$ for $m=3$ and $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}\geq\frac{5}{8}\texttt{CW}({\bf b}^{*})$ for $m=2$. These bounds are reproduced by substituting ${\rm tr}(H_{m}\rho)\geq\texttt{CW}({\bf b}^{*})$ and $\sum_{(j,k)\in E}w_{jk}\geq\texttt{CW}({\bf b}^{*})$ into Eq. (23).
@@
Because the cut weight is invariant to the bit flip, we can fix one of the $b_{j}$ to 00 or $1$. In this case, there exists a one-to-one relationship between $\mathcal{P}_{m}({\bf b};\rho^{*})+\mathcal{P}_{m}(\overline{\bf b};\rho^{*})$ and
${\mathcal{P}}_{m}(\mathfrak{p};\rho^{*},T)=\prod^{|V|-1}_{b=1}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{e_{<b}};\rho^{*},T),$ (25)
@@
is a random variable of an edge parity, where “$+$” and “--” indicate positive and negative parities, respectively, $T=\{e_{j_{b}}=(c_{j_{b}},p_{j_{b}})\}^{|V|-1}_{b=1}$ is an arbitrary spanning tree of given graph $G$, and $\mathfrak{p}_{e_{<b}}:=\{\mathfrak{p}_{e_{1}},\ldots,\mathfrak{p}_{e_{b-1}}\}$. Note that ${\bf b}^{\rm R}$ can be easily recovered using $T$ and $\mathfrak{p}^{\rm R}(T)$. Here, the reason for considering the probability of the edge parity on the tree is to exclude inconsistent parities such as a triangle with negative parities. Excluding such an inconsistency does not lose the generality because all possible configurations of solutions can be treated by considering only the probability on the tree. Using the edge parity expression, Eq. (20) can be rewritten as
$\begin{array}[]{ll}\mathfrak{p}^{\rm R}(T)&\displaystyle{=\mathop{\rm argmax}_{\mathfrak{p}\in\{+,-\}^{|V|-1}}\mathcal{P}_{m}(\mathfrak{p};\rho^{*},T)}\\ &\displaystyle{=\mathop{\rm argmax}_{\mathfrak{p}\in\{+,-\}^{|V|-1}}\prod^{|V|-1}_{b=1}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{e_{<b}};\rho^{*},T).}\end{array}$ (27)
In the following, we show that QRAO follows Eq. (20), whereas both the RQAOA and RQRAO follow Eq. (25) to obtain the solution. All of them assume a moderate approximation on a probability distribution. The following corollary is frequently used for the proof.
@@
Let us consider two assumptions: bit-flip and independent ones, which we refer to as assumptions A, denoted by the superscript A. The first one is the bit-flip assumption
${\mathop{{\rm argmax}}_{{\bf b}}\mathcal{P}^{\rm A}_{m}({\bf b};\rho^{*})}$
$=\mathop{{\rm argmax}}_{{\bf b}}\mathcal{P}^{\rm A}_{m}({\bf b};\rho^{*})+\mathcal{P}^{\rm A}_{m}(\overline{\bf b};\rho^{*}).$ (30)
In many cases, this assumption is practically satisfied for MAX-CUT . This is because the ground states of the MAX-CUT Hamiltonians are degenerate; ${\bf b}$ and $\overline{\bf b}$ represent the same cut. Moreover, unless a specific constraint is intentionally imposed to balance $\mathcal{P}_{m}({\bf b};\rho)$ and $\mathcal{P}_{m}(\overline{\bf b};\rho)$, optimization can result in one of $\mathcal{P}_{m}({\bf b};\rho)$ or $\mathcal{P}_{m}(\overline{\bf b};\rho)$ becoming large while the other becomes close to zero. Alternatively, in the case of an ansatz with $\mathbb{Z}_{2}$ symmetry [3], such as QAOA, the bit-flipped solution appears with the same probability, so the assumption is also satisfied in this case.
The second one is the independent assumption that all $b_{j_{a}}$ for $a=1,2,\ldots,|V|$ are independent where $\{j_{a}\}^{|V|}_{a=1}=\{1,...,|V|\}$, which is equivalent to write the joint distribution as
${\mathcal{P}}^{\rm A}_{m}({\bf b};\rho^{*})=\prod^{|V|}_{a=1}\mathcal{P}^{\rm A}_{m}(b_{j_{a}};\rho^{*}).$ (31)
Under these assumptions, Eq. (20) is approximated as
$\begin{array}[]{lll}b^{{\rm A}}_{j_{a}}&=&\displaystyle{\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\mathcal{P}^{{\rm A}}_{m}(b_{j_{a}};\rho^{*})}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{j_{a}}}}{2\sqrt{m}}\mathcal{E}_{j_{a}}\Bigg{)}}\\[14.22636pt] &=&\left\{\begin{array}[]{ll}0&(\mathcal{E}_{j_{a}}>0)\\ 1&(\mathcal{E}_{j_{a}}<0)\end{array}\right.,\end{array}$ (32)
where the second equality uses Corollary 3. Interestingly, ${\bf b}^{{\rm A}}$ is equivalent to the Pauli rounding of QRAO [11] defined in Eq. (14). This comes out that the Pauli rounding seeks the variable with the highest probability of ${\mathcal{P}}_{m}({\bf b};\rho)$ under the assumptions of Eqs. (30) and (31).
@@
Assume all $\mathfrak{p}_{e_{b}}$ for $b=1,2,\ldots,|V|-1$ are independent, which we refer to as assumption B, denoted by the superscript B. Under this assumption, Eq. (25) is approximated as
${\mathcal{P}}^{{\rm B}}_{m}(\mathfrak{p};\rho^{*},T)=\displaystyle{\prod_{b=1}^{|V|-1}\mathcal{P}^{{\rm B}}_{m}(\mathfrak{p}_{e_{b}};\rho^{*},T)},\$ (33)
and Eq. (27) is approximated as
$\begin{array}[]{lll}\mathfrak{p}^{{\rm B}}_{e_{b}=(c,p)}(T)&=&\displaystyle{\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}\in\{+,-\}}\mathcal{P}^{{\rm B}}_{m}(\mathfrak{p}_{e_{b}};\rho^{*},T)}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{(b_{c},b_{p})\in\{0,1\}^{2}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{c}+b_{p}}}{2m}\mathcal{E}_{cp}\Bigg{)}}\\[14.22636pt] &=&\left\{\begin{array}[]{ll}+&(\mathcal{E}_{cp}>0)\\ -&(\mathcal{E}_{cp}<0)\end{array}\right.,\end{array}$ (34)
where the second equality uses Corollary 3. We refer to this rounding method as tree rounding.
RQRAO with $N=1$, which has no recursive steps, is equivalent to using tree rounding. When $N=1$, the ensemble edge energy $\mathfrak{E}_{jk}$ in Eq. (18) is equal to the edge energy estimate $\mathcal{E}_{jk}$ in Eq. (17) from a single evaluation, and almost surely it takes a non-zero value. Since the spanning tree is constructed by gathering edges with $\mathfrak{E}_{jk}\neq 0$, a spanning tree covering all nodes can be formed with high probability. By progressively removing nodes from the leaves to the root of this spanning tree as illustrated in Fig. 4, all nodes except the root are eliminated, determining all parities at once. This is the same rounding method referred to as the tree rounding defined above. Since both Pauli rounding and tree rounding assume that all random variables are independent, it is expected that QRAO and RQRAO with $N=1$ will yield similar cut weights.
@@
Let us impose the following two assumptions, which we refer to as assumptions C, denoted by the superscript C:
${\max_{\mathfrak{p}\in\{+,-\}^{|V|-1}}\prod^{|V|-1}_{b=1}\mathcal{P}^{\rm C}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{<e_{b}};\rho^{*},T)}$
$=\prod^{|V|-1}_{b=1}\max_{\mathfrak{p}_{e_{b}}\in\{+,-\}}\mathcal{P}^{\rm C}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{<e_{b}};\rho^{*},T),$ (35)
$\mathcal{P}^{\rm C}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{<e_{b}};\rho^{*},T)=\mathcal{P}^{\rm C}_{m}(\mathfrak{p}_{e_{b}};\rho^{*}_{b}(\mathfrak{p}_{<e_{b}}),T),$ (36)
where $\rho^{*}_{b}(\mathfrak{p}_{<e_{b}})$ is the optimized quantum state that involves the information of $\mathfrak{p}_{<e_{b}}$. Under these assumptions, Eq. (27) can be approximated as
${\mathfrak{p}^{\rm C}_{e_{b}}(T)=\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}}\mathcal{P}^{{\rm C}}_{m}(\mathfrak{p}_{e_{b}};\rho^{*}_{b}(\mathfrak{p}_{<e_{b}}))}$
$\ \ \ {\rm for}\ \ \ b=1,\ldots,|V|-1.$ (37)
Compared to the Pauli and tree roundings, which impose the strong approximation on $\mathcal{P}_{m}$ that all random variables are independent, assumptions in Eqs. (36) and (37) are weak because each random variable is no longer independent. Although the assumption of independence in Pauli rounding and tree rounding was made due to the difficulty of conditional sampling, this approach addresses the problem by imposing conditions $\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}$ on the quantum state itself instead of on $\mathcal{P}_{m}$, where $\mathfrak{p}^{*}_{<e_{b}}$ is the determined edge parities before step $b$. Of course, this approach is not exactly equivalent to the original one, but imposing the condition as a constraint in the optimization should provide a good approximation to conditional sampling. In the following, we specifically demonstrate how the method of recursively decoding bit strings, as adopted in Recursive QAOA [3], follows Eq. (27) to obtain ${\mathfrak{p}^{\rm R}(T)}$ under the assumptions in Eqs. (36) and (37).
@@
Finally, $\mathfrak{p}^{*}=\{\mathfrak{p}^{*}_{1},\ldots,\mathfrak{p}^{*}_{|V|-1}\}$ is obtained. The corresponding bit string ${\bf b}^{\rm C}$ can be easily recovered from $\mathfrak{p}^{*}$ by fixing one of the $b^{\rm C}_{j}$ to 00 or $1$. The condition $\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}$ is indirectly imposed through $\rho^{*}_{b}$ to $\mathcal{P}_{m}$, and hence this rounding method imposes an approximation that is weaker than that of Pauli and tree rounding. We refer to this rounding method as recursive rounding.